# Xinlu Zhang1, Chenxin Tian2, Xianjun Yang1, Lichang Chen3, Zekun Li1, Linda Ruth Petzold 1

### v4

@@ 6 Human Study
Figure 3: Human study results. Head-to-head clinician preference comparison between AlpaCare-13B and PMC-13B on 100 medical free-form instruction instances on (a) correctness and (b) helpfulness.
We further conduct human studies to label question-and-answer pairs in medical free-form instruction evaluation. Three annotators with MD degrees in progress are involved in the study to perform pairwise comparisons for each question and answer pair. Specifically, we randomly select 50 prompts from each test set, totaling 100 prompts. These prompts, along with the responses generated by both AlpaCare-13B and PMC-13B, the best baseline in the 13B models, are presented to the annotators for evaluation. The evaluation is based on two criteria: correctness and helpfulness. Correctness evaluates whether the response provides accurate medical knowledge to address the question posed, while helpfulness measures the model’s ability to assist users concisely and efficiently, considering the user intent. In practical terms, an answer can be correct but not necessarily helpful if it is too verbose and lacks guidance. To determine the final result for each criterion of each evaluation instance, we employ a majority vote method. If at least two of the annotators share the same opinion, their preference is considered the final answer; otherwise, we consider the outputs of the two models to be tied. The results are shown in Figure 3.
Consistent with previous results, the AlpaCare-13B obtains better performance than PMC-13B in the human evaluation. The output generated by AlpaCare-13B is more accurate than that from PMC-13B, with 54% answers preferred by expert annotators. This result shows a superior medical capacity of AlpaCare. Beyond correctness, AlpaCare generates answers that are significantly more helpful compared to PMC-13B. 69% answers generated by AlpaCare are preferred by domain experts, highly exceeding the 11% preference rate for PMC-13B outputs, demonstrating better practical usability. The higher relative gain in helpfulness of AlpaCare compared to correctness is expected because the goal of IFT is to obtain the instruction-following ability of LLMs to align the model with various user needs, rather than gaining new knowledge (Zhou et al., 2023).
@@ 7 Analysis & Case Study
7.1 IFT Dataset Diversity Analysis
Training a model with diverse instructions enhances its ability to follow instructions (Wang et al., 2023b; Chen et al., 2023). However, current medical LLMs often have training data lacking in instructional diversity, typically using repetitive instructions across different instances (Li et al., 2023b; Han et al., 2023; Wu et al., 2023). To examine the diversity in our dataset, we plot the distributions of 4 key areas for instruction generation from MedInstruct-52k, shown in Figure 4 (a)- (d). Specifically, we present the top 20 topics, views and types and the difficulty levels from 1 to 5, offering insight into training data distribution. We further analyze instruction linguistic diversity by showing the root verbs and their corresponding direct-object nouns from each instruction. The top 20 root verbs and their 4 most common direct-object nouns are displayed in Figure 4 (e), representing 22% of the total dataset. Our findings show quite diverse medical intents and textual formats in our MedInstruct-52k.
@@ 7.1 IFT Dataset Diversity Analysis
To quantitatively showcase our dataset’s diversity in comparison to IFT dataset of other medical LLMs, we calculate the linguistic entropy in the instructions of instruction-following datasets used for medical IFT models. Higher entropy values signify greater diversity. Specifically, we analyze the top 20 root verbs and their 4 primary direct noun objects for each dataset and calculate verb-noun pair entropy, as shown in Table 8. AlpaCare’s dataset, MedInstruct-52K, exhibits the highest entropy, underscoring its superior diversity, which in turn better elicits the model’s instruction-following capabilities in the IFT stage.
@@
7.2 Generation Case Study
Figure 5(a) showcases a case on correctness with a high medical difficulty level. LLAMA-13B struggles to provide correct answers to the questions in the instruction. In contrast, Medical LLMs demonstrate a marked improvement in the response accuracy. PMC offers a general overview, citing common causes like stress, hormonal imbalances, and nutritional deficiencies. However, it falls short by using ambiguous terms like "certain medical conditions" and "an underlying medical condition," which lack specificity and fail to deliver actionable medical insights. Conversely, AlpaCare provides a detailed analysis, pointing specific conditions such as hypothyroidism and iron deficiency anemia that could explain the patient’s symptoms and offering a nuanced understanding that significantly enhances the precision and utility of the guidance provided. Additionally, it emphasizes the importance of medical attention for severe symptoms, making its recommendations more actionable.
In the case of helpfulness illustrated in Figure 5(b), both the base model and medical models generate accurate medical information on Metoprolol from different perspectives. However, LLAMA and PMC do not adhere to the instruction to use a bullet-point format, which would enhance clarity. Conversely, AlpaCare clearly follows the instruction with well-structured formatting, demonstrating that fine-tuning with a diverse instruction-following dataset improves the ability to follow instructions and, in turn, increases helpfulness.
@@ 8 Conclusion
In conclusion, our study fills a gap in aligning LLMs with domain-specific user intents in the medical domain while managing costs. Specifically, we propose a semi-automated pipeline using GPT-4 and ChatGPT with a high-quality expert-curated seed set to create a diverse machine-generated medical IFT dataset, MedInstruct-52k, for LLM tuning. The resulting model, AlpaCare, demonstrates strong medical instruction-following and robust generalization abilities compared to previous medical LLMs. Our approach shows versatility across various metrics, including different LLM judges (such as ChatGPT and Claude-2), a range of base model families (like LLaMA-1 and LLaMA-2), and various model sizes (7B and 13B). This highlights the significant benefits of incorporating diverse data in medical AI model development.
9 Ethics Statement
This research demonstrates the potential of enhancing open-source LMs’ medical capacity and general applicability by distilling knowledge from more powerful "teacher" LLMs. We improve the smaller models’ ability to follow medical instructions and align with user intentions, offering potential benefits for future healthcare applications. Even though our model AlpaCare maintains comparable truthfulness scores compared to other baselines, but the rate of correct answers is still low, showing that our model is likely to hallucinate incorrect answers. Therefore, it is crucial to emphasize that this system is designed to serve as an assistant tool, complementing but not replacing the expertise and judgment of healthcare professionals. All outputs generated by this system must be rigorously validated by licensed medical practitioners before any practical application.
The transition of these LMs to practical medical scenarios — useful for healthcare professionals, patients, and other medical personnel — necessitates extensive research to ensure safety, privacy, and reliability. This development involves not only technical robustness but also a commitment to ethical standards. It includes comprehensive quality assessments in various clinical environments to ensure that the system meets the highest standards of accuracy and ethical conduct. Additionally, attention must be paid to the potential implications for patient privacy and data security, ensuring that all patient information is handled with the utmost confidentiality and in compliance with relevant data protection laws and healthcare regulations.
Acknowledgments
We gratefully acknowledge the generous financial support provided by the National Institutes for Health (NIH) grant NIH 7R01HL149670.
Appendix A Appendix
Limitations
Our approach utilizes ’teacher’ LLMs, such as GPT-4 and ChatGPT, to automatically generate medical instruction-response pair datasets, employing these teacher models as medical knowledge bases. However, this could result in hallucinations in the medical knowledge generation. To enhance the generation reliability, we aim to integrate LLMs with the internet and knowledge graphs in future work.
A.1 Medical Task Difficulty Level Scoring System
@@
We introduce a clinician-crafted seed set to generate MedInstruct-52k and a free-form medical instruction evaluation set, MedInstruct-test . This set spans a medical difficulty scale ranging from 1 to 5, where 1 represents the easiest tasks and 5 indicates the most challenging ones. A clinician assessed the difficulty levels of all instances within both the seed set and MedInstruct-test based on the scoring system shown in Table 9 This system offers a refined dimension for prompting GPT-4 to produce tasks across varied difficulty levels and to evaluate medical proficiency of IFT models.
A.2 Prompt details for MedInstruct-52k generation
Here we provides prompts we use for query GPT-4 and ChatGPT for task and response generation.
@@ A.2 Prompt details for MedInstruct-52k generation
A.3 Training hyperparameter details
@@
A.4 More experimental results
A.4.1 General domain free-form instruction evaluation
Table 13: Comparison on general domain free-form instruction evaluation. A performance comparison between AlpaCare and IT baselines on AlpacaFarm on 4 distinct reference models: Text-davinci-003, GPT-3.5-turbo, GPT-4 and Claude-2. ‘AVG’ represents the mean performance score across all referenced models.
@@ A.4.1 General domain free-form instruction evaluation
A.4.2 Ablation study
@@
iCliniq
Text-davinci-003 GPT-3.5-turbo GPT-4 Claude-2 AVG
LLaMA Alpaca 38.8 30.4 12.8 15.6 24.4
AlpaCare 66.6 50.6 47.4 49.7 53.6
LLaMA-2 Alpaca 45.8 36.3 18.2 20.8 30.3
AlpaCare 66.5 50.4 47.8 50 53.7
MedInstruct
Text-davinci-003 GPT-3.5-turbo GPT-4 Claude-2 AVG
LLaMA Alpaca 35.0 20.6 21.5 15.6 24.4
AlpaCare 67.6 48.8 47.4 49.7 53.5
LLaMA-2 Alpaca 39.6 22.7 26.4 18.5 26.8
AlpaCare 70.6 48.8 50.0 48.4 54.2
iCliniq
Text-davinci-003 GPT-3.5-turbo GPT-4 Claude-2 AVG
Alpaca 46.7 37.0 19.6 21.7 31.3
Medalpaca 8.1 4.4 1.0 2.0 3.9
PMC 40.6 29.0 14.3 17.5 25.4
AlpaCare 66.7 51.2 48.2 50.2 54.4
MedInstruct
Text-davinci-003 GPT-3.5-turbo GPT-4 Claude-2 AVG
Alpaca 39.8 22.5 27.1 18.1 26.9
Medalpaca 0.2 0 0 0 0.1
PMC 44.9 31.9 32.8 29.2 34.7
AlpaCare 71.3 49.1 49.8 47.7 54.5

### v6

@@ 6 Human Study
Figure 3: Human study results. Head-to-head clinician preference comparison between AlpaCare-13B and PMC-13B on (a) correctness and (b) helpfulness.
We further conduct human studies to label question-and-answer pairs in medical free-form instruction evaluation. Three annotators with MD degree in progress are involved in the study to perform pairwise comparisons for each question and answer pair. Specifically, we randomly select 50 prompts from each test set, totaling 100 prompts. These prompts, along with the responses generated by both AlpaCare-13B and PMC-13B, the best baseline in the 13B models, are presented to the annotators for evaluation. The evaluation is based on two criteria: correctness and helpfulness. Correctness evaluates whether the response provides accurate medical knowledge to address question posed, while helpfulness measures the model’s ability to assist users concisely and efficiently, considering the user intent. In practical terms, an answer can be correct but not necessarily helpful if it is too verbose and lacks guidance. To determine the final result for each criterion of each evaluation instance, we employ a majority vote method. If at least two of the annotators share the same opinion, their preference is considered the final answer; otherwise, we consider the outputs of the two models to be tied. The results are shown in Figure 3.
Consistent with previous results, AlpaCare-13B outperforms PMC-13B in human evaluation, with 54% of answers preferred by expert annotators for correctness and 69% for helpfulness. This demonstrates AlpaCare’s superior medical capacity and practical usability. The greater improvement in helpfulness over correctness for AlpaCare is expected, as the goal of IFT is to enhance LLMs’ instruction-following ability to meet diverse user needs, rather than acquiring new knowledge.
@@ 7 Analysis & Case Study
IFT Dataset Diversity Analysis
Training a model with diverse instructions enhances its ability to follow instructions (Wang et al., 2023b). However, current medical LLMs often have training data lacking in instructional diversity, typically using repetitive instructions across different instances (Li et al., 2023b; Han et al., 2023; Wu et al., 2023). To examine the diversity in our dataset, we plot the distributions of 4 key areas for instruction generation from MedInstruct-52k, shown in Figure 4 (a)- (d). Specifically, we present the top 20 topics, views and types and the difficulty levels from 1 to 5, offering insight into training data distribution. We further analyze instruction linguistic diversity by showing the root verbs and their corresponding direct-object nouns from each instruction. The top 20 root verbs and their 4 most common direct-object nouns are displayed in Figure 4 (e), representing 22% of the total dataset. Our findings show quite diverse medical intents and textual formats in our MedInstruct-52k.
@@
To quantitatively showcase our dataset’s diversity in comparison to IFT dataset of other medical LLMs, we calculate the linguistic entropy in the instructions of instruction-following datasets used for medical models. Higher entropy values signify greater diversity. Specifically, we analyze the top 20 root verbs and their 4 primary direct noun objects for each dataset and calculate verb-noun pair entropy, as shown in Table 7. AlpaCare’s dataset, MedInstruct-52K, exhibits the highest entropy, underscoring its superior diversity, which in turn better elicits the model’s instruction-following capabilities during IFT.
Generation Case Study
Figure 5(a) illustrates a correctness case with a high medical difficulty level, where LLAMA-13B struggles to provide accurate responses, while Medical LLMs show improvement. However, PMC provides a general overview, mentioning common causes like stress and hormonal imbalances, and uses vague terms such as ’certain medical conditions,’ which lack specificity and do not provide actionable medical insights. In contrast, AlpaCare gives a detailed analysis, identifying specific conditions like hypothyroidism and iron deficiency anemia, and emphasizes the importance of medical attention for severe symptoms, enhancing the guidance’s precision and actionability.
Figure 5(b) shows a helpfulness case study. Both the base and medical models accurately describe Metoprolol, but LLAMA and PMC disregard the requested bullet-point format, reducing clarity. In contrast, AlpaCare effectively follows the instructions with well-organized formatting. This demonstrates that fine-tuning with a diverse medical IFT dataset enhances the model’s ability to follow instructions, thereby increasing helpfulness.
@@ 8 Conclusion
In this paper, we propose a semi-automated pipeline using GPT-4 and ChatGPT to create the diverse MedInstruct-52k dataset for LLM tuning. Extensive experiments on multiple benchmarks show that when trained on this dataset, our model, AlpaCare, demonstrates stronger medical and general instruction-following capabilities compared to medical LLM baselines, underscoring the importance of data diversity in medical AI model development.
Appendix A Data and Code
We provide codebase, MedInstruct-52k and MedInstruct-test, and AlpaCare models in the link https://anonymous.4open.science/r/AlpaCare-D6BB/.
Appendix B Medical Task Difficulty Level Scoring System
We introduce a clinician-crafted seed set to generate MedInstruct-52k and a free-form medical instruction evaluation set, MedInstruct-test. This set spans a medical difficulty scale ranging from 1 to 5, where 1 represents the easiest tasks and 5 indicates the most challenging ones. A clinician assessed the difficulty levels of all instances within both the seed set and MedInstruct-test based on the scoring system shown in Table 8 This system offers a refined dimension for prompting GPT-4 to produce tasks across varied difficulty levels and to evaluate medical proficiency of instruction-tuned models.
@@ Appendix B Medical Task Difficulty Level Scoring System
Appendix C Prompt details for MedInstruct-52k generation
Here we provide prompts we use for query GPT-4 and ChatGPT for task and response generation.
@@ Appendix C Prompt details for MedInstruct-52k generation
Appendix D LLM APIs cost analysis and selection reasons
We spent approximately \$900 on task generation using GPT-4 and \$500 on response generation using GPT-3.5-Turbo. The cost of using the GPT-4 API is 30 times higher than that of GPT-3.5-Turbo at the time we created the dataset, which would increase the response cost to \$15,000, making it difficult to stay within budget. Therefore, we utilized GPT-3.5-Turbo as the response generator and one of our evaluators to balance cost and effectiveness. To demonstrate the suitability of using GPT-3.5-Turbo for dataset generation, we verified the quality of MedInstruct-52k by randomly selecting 50 instances and having a clinician evaluate the responses generated by GPT-3.5-Turbo. The clinician found 49 out of 50 responses to be correct, demonstrating the dataset’s high quality.
Appendix E Training hyperparameter details
Here we provide hyperparameter details for AlpaCare training.
@@ Appendix E Training hyperparameter details
Appendix F MedInstruct-test statistics
Here, we present the test dataset statistics in Table 12.
Difficulty Level 1 2 3 4 5
Count 44 46 41 41 44
Appendix G More experimental results
General domain free-form instruction evaluation
Table 13: Comparison on general domain free-form instruction evaluation. A performance comparison between AlpaCare and baselines on AlpacaFarm on 4 distinct reference models: Text-davinci-003, GPT-3.5-turbo, GPT-4 and Claude-2. ‘AVG’ represents the mean performance score across all referenced models.
@@ Appendix G More experimental results
More analysis in general domain performance
Model StrategyQA DROP
Alpaca 57.80 23.68
AlpaCare 58.02 24.96
Table 14: Performance of Alpaca and AlpaCare on StrategyQA and DROP datasets.
Compared to Alpaca (Taori et al., 2023), AlpaCare achieves better results in the general domain. This improvement is likely due to the intensive knowledge and reasoning embedded in the medical dataset (Liévin et al., 2023; Chen et al., 2024). For example, in the BBH results, the top three categories where AlpaCare outperforms Alpaca are ‘dyck_languages’, ‘movie_recommendation’, and ‘navigate’, which requires strong knowledge and reasoning abilities. To further support these findings, we conducted a knowledge-intensive commonsense evaluation using StrategyQA (Geva et al., 2021) and an additional reasoning benchmark evaluation using DROP (Dua et al., 2019) to compare Alpaca and AlpaCare with LLaMA (Touvron et al., 2023a)-7B as backbone, following the methodologies in Gao et al. (2021) and Chia et al. (2023), respectively. The Table 14 presents the results.
These results reinforce that AlpaCare’s enhanced performance is not limited to the medical domain but also extends to broader general domain tasks, thereby confirming its superior generalizability.
Ablation study
We show the detailed score of 4 reference models for medical free-form instruction evaluation on 13B instruction-tuned models in Table 15.
iCliniq
Text-davinci-003 GPT-3.5-turbo GPT-4 Claude-2 AVG
Alpaca 46.7 37.0 19.6 21.7 31.3
Medalpaca 8.1 4.4 1.0 2.0 3.9
PMC 40.6 29.0 14.3 17.5 25.4
AlpaCare 66.7 51.2 48.2 50.2 54.4
MedInstruct
Text-davinci-003 GPT-3.5-turbo GPT-4 Claude-2 AVG
Alpaca 39.8 22.5 27.1 18.1 26.9
Medalpaca 0.2 0 0 0 0.1
PMC 44.9 31.9 32.8 29.2 34.7
AlpaCare 71.3 49.1 49.8 47.7 54.5
We show the detailed score of 4 reference models for medical free-form instruction evaluation on different backbones in Table 16.
iCliniq
Text-davinci-003 GPT-3.5-turbo GPT-4 Claude-2 AVG
LLaMA Alpaca 38.8 30.4 12.8 15.6 24.4
AlpaCare 66.6 50.6 47.4 49.7 53.6
LLaMA-2 Alpaca 45.8 36.3 18.2 20.8 30.3
AlpaCare 66.5 50.4 47.8 50 53.7
LLaMA-3 Alpaca 42.3 28.6 26.4 10.0 26.8
AlpaCare 77.6 53.9 46.3 49.7 56.9
MedInstruct
Text-davinci-003 GPT-3.5-turbo GPT-4 Claude-2 AVG
LLaMA Alpaca 35.0 20.6 21.5 15.6 24.4
AlpaCare 67.6 48.8 47.4 49.7 53.5
LLaMA-2 Alpaca 39.6 22.7 26.4 18.5 26.8
AlpaCare 70.6 48.8 50.0 48.4 54.2
LLaMA-3 Alpaca 38.4 16.9 14.6 13.0 20.7
AlpaCare 78.5 50.0 51.4 46.5 56.6
We show the detailed score of 4 reference models for medical free-form instruction evaluation by using Claude-2 as the judge in Table 17.